OpenStudy (anonymous):

Can someone show/tell me step by step on how to change y-5=1/8(x-11) Into standard form? (ax+by=c) (the 1/8 means one-eighth) Thanks!!

OpenStudy (imstuck):

First of all you need to do all the math to get it into slope-intercept form. combine all the like terms and get a y = equation.

OpenStudy (imstuck):

\[y-5=\frac{ 1 }{8 }x-\frac{ 11 }{8 }\]That's the first step. Next step is\[y=\frac{ 1 }{8 }x-\frac{ 11 }{ 8 }+5\]Next step is\[y=\frac{ 1 }{8 }x-\frac{ 11 }{8 }+\frac{ 40 }{ 8 }\]

OpenStudy (imstuck):

Next is to clean it up completely. Like this:\[y=\frac{ 1 }{ 8 }x+\frac{ 29 }{ 8 }\]Finally to get it into the form you want it, which is standard form, you move the x over to the same side as the y and have it equal your constant. Like this:

OpenStudy (imstuck):

\[y-\frac{ 1 }{ 8 }x=\frac{ 29 }{ 8 }\]Most books and teachers do not like the fractions, so to get rid of the 8 in the denominators, you will multiply every term by 8 and you get this:\[8y-x=29\]

OpenStudy (imstuck):

That's the equation in standard form. Hope the step by step helped.

OpenStudy (anonymous):

It helped so much!! Thank you!

OpenStudy (precal):

standard form is Ax+By=C

OpenStudy (precal):

A should be a positive value and a whole number